### v1

# Compressed Image Generation with Denoising Diffusion Codebook Models

Abstract
We present a novel generative approach based on Denoising Diffusion Models (DDMs), which produces high-quality image samples along with their losslessly compressed bit-stream representations. This is obtained by replacing the standard Gaussian noise sampling in the reverse diffusion with a selection of noise samples from pre-defined codebooks of fixed iid Gaussian vectors. Surprisingly, we find that our method, termed Denoising Diffusion Codebook Model (DDCM), retains sample quality and diversity of standard DDMs, even for extremely small codebooks. We leverage DDCM and pick the noises from the codebooks that best match a given image, converting our generative model into a highly effective lossy image codec achieving state-of-the-art perceptual image compression results. More generally, by setting other noise selections rules, we extend our compression method to any conditional image generation task (e.g., image restoration), where the generated images are produced jointly with their condensed bit-stream representations. Our work is accompanied by a mathematical interpretation of the proposed compressed conditional generation schemes, establishing a connection with score-based approximations of posterior samplers for the tasks considered. Code and demo are available on our project’s website.
Machine Learning, ICML
Figure 1: Our proposed scheme (DDCM) produces visually appealing image samples with high compression ratios (bottom-right corners).
1 Introduction
Denoising Diffusion Models (DDMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020) have emerged as an effective tool for generating samples from complex signal distributions (e.g., natural images). Hence, DDMs are commonly leveraged to facilitate a variety of downstream tasks, such as text-to-image synthesis (Ramesh et al., 2021; Rombach et al., 2022; Saharia et al., 2022), editing (Meng et al., 2022; Huberman-Spiegelglas et al., 2024), compression (Theis et al., 2022; Elata et al., 2024; Körber et al., 2024), and restoration (Kawar et al., 2022; Chung et al., 2023). Common to many of these applications is the reliance on iterative sampling from a continuous Gaussian distribution, yielding an unbounded representation space.
This work embarks on the hypothesis that such an infinite representation space is highly redundant. For example, consider any stochastic diffusion generative process with $T=1000$ sampling steps (e.g., DDPM (Ho et al., 2020)). Suppose that at each timestep, the generative process is restricted to choosing between only two fixed noise realizations. Sampling could then lead to $2^{1000}$ different outputs, an incredibly large number exceeding the estimated amount of atoms in the universe. Thus, in principle, such a process could cover the distribution of natural images densely.
We harness this intuition and propose Denoising Diffusion Codebook Models (DDCM), a novel DDM generation scheme for continuous signals, leveraging a discrete and finite representation space. In particular, we first construct a chain of codebooks, where each is a sequence of pre-sampled Gaussian noise vectors. These codebooks are constructed once and remain fixed for the entire lifetime of the model. Then, during the generative process, we simply randomly pick the noises from the codebooks instead of drawing them from a Gaussian distribution, as shown in Fig. 2. Since we alter only the sampling process, DDCM can be applied using any pre-trained DDM. Interestingly, we find that our proposed discrete and finite representation space is indeed expressive enough to retain the generative capabilities of standard DDMs, even when using incredibly small codebooks. Since our generative process is entirely governed by the noise indices picked during the generation, an important consequence is that every generated image can be perfectly reconstructed by repeating the process with its corresponding indices.leverage this property to solve a variety of tasks, using gradient-free noise selection rules to guide the DDCM generation process. In particular, by choosing the discrete noises to best match a given image, we achieve state-of-the-art perceptual compression results. Moreover, using DDCM with different noise selection rules yields a versatile framework for other compressed conditional generation tasks, such as compressed image restoration (see examples in Fig. 1). Finally, we provide a mathematical interpretation of the proposed schemes based on score-based generative modeling with SDEs (Song et al., 2020), showing a connection between our generalized selection rules and approximate posterior sampling for compressed conditional generation.
Figure 2: Method overview. DDCM replaces the standard Gaussian noises in DDPM sampling with a selection of noise samples from pre-defined codebooks of fixed iid Gaussian vectors. This retains the high-quality generative properties of standard DDMs, while producing the results along with their compressed representations. By choosing the discrete noises according to different selection rules, DDCM can perform a variety of conditional image generation tasks. Our highly condensed bit-stream representation is especially effective for image compression, leading to state-of-the-art results.
2 Related Work
Compression.
Image compression has seen significant progress in recent years, with the penetration of neural networks to this domain. Neural methods range from constructing specialized architectures (Ballé et al., 2017; Zhu et al., 2022; Jiang et al., 2023; Jiang & Wang, 2023) to relying on different generative models such as GANs (Mentzer et al., 2020; Muckley et al., 2023; Iwai et al., 2024) and VAEs (Theis et al., 2017). Recent compression methods leverage DDMs and offer high perceptual quality results, by training models from scratch (Yang & Mandt, 2023; Ghouse et al., 2023), fine-tuning existing models (Careil et al., 2023; Körber et al., 2024), or using pre-trained DDMs in a zero-shot manner (without further training) (Theis et al., 2022; Elata et al., 2024). Current solutions in the latter category are highly computationally demanding, either due to their communication schemes (e.g., reverse channel coding (Theis & Ahmed, 2022; Theis et al., 2022)) or their need to perform thousands of denoising operations (Elata et al., 2024). Our work falls into this last category, offering a novel and highly effective compression scheme with a fast bit-stream communication method and computational demands that match standard use of DDMs.
Discrete Generative Modeling for Continuous Data.
Recent works have explored discrete generative modeling of continuous data distributions. These employ various discrete representations, such as vector quantized latent tokens (Wu et al., 2024) or hierarchical modeling schemes that gradually refine each generated sample (Yang, 2024). DDCM offers a new such discrete generative framework, building on the exceptional achievements of DDMs.
Conditional Image Generation with Pre-Trained DDMs.
Pre-trained DDMs are commonly utilized for solving conditional image generation tasks, such as image restoration (Kawar et al., 2022; Lugmayr et al., 2022; Wang et al., 2023b; Chung et al., 2023; Lin et al., 2024; Song et al., 2023; Cohen et al., 2024b; Yue & Loy, 2024; Raphaeli et al., 2025; Man et al., 2025) and editing (Meng et al., 2022; Huberman-Spiegelglas et al., 2024; Cohen et al., 2024a; Manor & Michaeli, 2024). In this work we address these tasks from the lens of DDCM, where the conditional samples are generated along with their compressed bit-streams.
Compressed Image Generation.
The task of compressed image generation (generating images directly in their compressed form) has been previously explored. Kang et al. (2019) trained an unconditional Generative Adversarial Network (GAN) (Goodfellow et al., 2014) to synthesize JPEG representations. Rajesh et al. (2023) trained a text-conditional GAN in the JPEG domain. In this work, we use DDMs instead of GANs and introduce a novel compressed image representation space different than that of JPEG. Our approach is compatible with any pre-trained DDM without requiring additional training.
3 Background
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020) generate samples from a data distribution $p_{0}$ by reversing a diffusion process that gradually adds random noise to samples from the data. Specifically, the diffusion process starts with ${\mathbf{x}}_{0}\sim p_{0}$ and produces the chain ${\mathbf{x}}_{0},{\mathbf{x}}_{1},~{}\ldots~{},{\mathbf{x}}_{T}$ via
${\mathbf{x}}_{i}=\sqrt{\alpha_{i}}{\mathbf{x}}_{i-1}+\sqrt{1-\alpha_{i}}{\mathbf{z}}_{i},\quad{\mathbf{z}}_{i}\sim\mathcal{N}({\bm{0}},{\bm{I}}),$ (1)
where $\alpha_{1},\ldots,\alpha_{T}>0$ are some time-dependent constants. The above is a time-discretization of a Variance Preserving (VP) SDE (Song et al., 2020). Then, samples from the data distribution $p_{0}$ are generated by solving the corresponding reverse-time VP SDE (Anderson, 1982; Song et al., 2020), i.e., by gradually denoising samples, starting from ${\mathbf{x}}_{T}\sim\mathcal{N}({\bm{0}},{\bm{I}})$. In this paper we adopt Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020), which propose the generative process ($i=T,\ldots,1$)
${\mathbf{x}}_{i-1}={\bm{\mu}}_{i}({\mathbf{x}}_{i})+\sigma_{i}{\mathbf{z}}_{i},~{}\mbox{where}$ (2)
${\bm{\mu}}_{i}({\mathbf{x}}_{i})=\frac{1}{\sqrt{\alpha_{i}}}({\mathbf{x}}_{i}+(1-\alpha_{i}){\bm{s}}_{i}({\mathbf{x}}_{i})),$ (3)
${\mathbf{z}}_{i}\sim\mathcal{N}({\bm{0}},{\bm{I}})$, $\sigma_{i}=\sqrt{1-\alpha_{i}}$, and ${\bm{s}}_{i}({\mathbf{x}}_{i})$ denotes the score $\nabla_{{\mathbf{x}}_{i}}\log{p_{i}({\mathbf{x}}_{i})}$ of the probability density function $p_{i}({\mathbf{x}}_{i})$. Such a score function ${\bm{s}}_{i}({\mathbf{x}}_{i})$ is typically learned via denoising score matching (Vincent, 2011; Song & Ermon, 2019; Song et al., 2020; Ho et al., 2020), where a model $\hat{{\mathbf{x}}}_{0|i}$ is trained to predict ${\mathbf{x}}_{0}$ from ${\mathbf{x}}_{i}$ (i.e., a denoiser), and using the well-known equation (Robbins, 1956; Miyasawa et al., 1961; Stein, 1981)
${\bm{s}}_{i}({\mathbf{x}}_{i})=\frac{\sqrt{\bar{\alpha}_{i}}\hat{{\mathbf{x}}}_{0|i}-{\mathbf{x}}_{i}}{1-\bar{\alpha}_{i}},$ (4)
where $\bar{\alpha}_{i}\coloneqq\prod_{s=1}^{i}\alpha_{s}$. This generative process is applicable to both pixel space (Dhariwal & Nichol, 2021) and latent space (Rombach et al., 2022) diffusion models, by employing a VAE-based encoder-decoder.
4 Denoising Diffusion Codebook Models
Method.
Equation (2) depicts the standard DDPM sampling approach, where the added noise is sampled from a continuous Gaussian distribution. DDCM instead uses a discrete noise space, by limiting each sampling step to choose from $K$ constant noise realizations, fixed separately for each step. Formally, for each $i=2,\ldots,T+1$ we define a codebook of $K$ entries
${\mathcal{C}}_{i}=\left[{\bm{z}}_{i}^{(1)},{\bm{z}}_{i}^{(2)},\ldots,{\bm{z}}_{i}^{(K)}\right],$ (5)
where each ${\mathcal{C}}_{i}(k)\coloneqq{\bm{z}}_{i}^{(k)}$ is sampled independently from $\mathcal{N}({\bm{0}},{\bm{I}})$ and remains fixed throughout the model’s lifetime. Then, we modify the DDPM sampling process (2), replacing the noise ${\mathbf{z}}_{i}$ by a randomly selected codebook entry,
${\mathbf{x}}_{i-1}={\bm{\mu}}_{i}({\mathbf{x}}_{i})+\sigma_{i}{\mathcal{C}}_{i}(k_{i}),$ (6)
where the process is initialized with ${\mathbf{x}}_{T}={\mathcal{C}}_{T+1}(k_{T+1})$, $k_{i}\sim\text{Unif}(\{1,\ldots,K\})$, and sampling step $i=1$ does not involve noise addition. This random selection procedure is depicted in the generation branch in Fig. 2. Importantly, running the generative process (6) with a given sequence of noise vectors $\{{\mathcal{C}}_{i}(k_{i})\}_{i=2}^{T+1}$ always produces the same output image. Thus, as depicted in the bottom part of Fig. 2, the sequence of indices $k_{T+1},\ldots,k_{2}$ can be considered a losslessly compressed bit-stream representation of each generated image.
Figure 3: Comparing DDPM with DDCM for different codebook sizes $K$. Interestingly, DDCM with $K=64$ achieves similar FID to DDPM, suggesting that the continuous representation space of DDPM (DDCM with $K=\infty$) is highly redundant. We use a class-conditional ImageNet model ($256\times 256$) for pixel space, and the text-conditional SD 2.1 model ($768\times 768$) for latent space, with prompts from MS-COCO. The $K$ axis is in log-scale.
Experiments.
While DDPM is equivalent to DDCM with $K=\infty$, the first question we address is whether DDCM maintains the synthesis capabilities of DDPM for relatively small $K$ values. We compare the performance of DDPM with that of DDCM using $K\in\{2,4,8,16,64\}$ for sampling from pre-trained pixel and latent space models. We compute the Fréchet Inception Distance (FID) (Heusel et al., 2017) to evaluate the generation performance. In LABEL:app:more_random_gens we report additional metrics and provide qualitative comparisons. For pixel space generation, we use a class-conditional DDM trained on ImageNet $256\times 256$ (Deng et al., 2009; Dhariwal & Nichol, 2021), and apply classifier guidance (CG) (Dhariwal & Nichol, 2021) with unit scale. We use the 50k validation set of ImageNet as the reference dataset, and sample 10k class labels randomly to generate the images. For latent space, we use Stable Diffusion (SD) 2.1 (Rombach et al., 2022) trained on $768\times 768$ images and apply classifier-free guidance (CFG) with scale 3 (equivalent to $w=2$ in (Ho & Salimans, 2021)). As the reference dataset, we randomly select 10k images from MS-COCO (Lin et al., 2014; Chen et al., 2015) along with one caption per image, and use those captions as prompts for sampling.
As shown in Fig. 3, DDCM achieves similar FID scores to DDPM at $K=64$, suggesting that the Gaussian representation space of DDPM is redundant. In the next sections we leverage our new representation space to solve a variety of tasks, including image compression and compressed image restoration.
5 Image Compression with DDCM
Figure 4: Qualitative image compression results. The presented images are taken from the Kodak24 ($512\times 512$) dataset. Our codec produces highly realistic outputs, while maintaining better fidelity to the original images compared to previous methods.
Method.
Since sampling with DDCM yields compact bit-stream representations, a natural endeavor is to harness DDCM for compressing real images. In particular, to compress an image ${\mathbf{x}}_{0}$, we leverage the predicted $\hat{{\mathbf{x}}}_{0|i}$ (Eq. 4) at each timestep $i$ and compute the residual error from the target image, ${\mathbf{x}}_{0}-\hat{{\mathbf{x}}}_{0|i}$. Then, we guide the sampling process towards ${\mathbf{x}}_{0}$ by selecting the codebook entry that maximizes the inner product with this residual,
$k_{i}=\operatorname*{arg\,max}_{k\in\{1,\ldots,K\}}\langle{\mathcal{C}}_{i}(k),{\mathbf{x}}_{0}-\hat{{\mathbf{x}}}_{0|i}\rangle,$ (7)
where the size of the first codebook ${\mathcal{C}}_{T+1}$ is $K=1$. This process is depicted as the compression branch in Fig. 2, where the resulting set of chosen indices $\{k_{i}\}_{i=2}^{T+1}$ is the compressed bit-stream representation of the given image. Section 6 sheds more light on this choice of the noise selection from the perspective of score-based generative models (Song et al., 2020). As in Sec. 4, decompression follows standard DDCM sampling 6, re-selecting the stored indices instead of picking them randomly. When using latent space DDMs (e.g., SD), we first encode ${\mathbf{x}}_{0}$ into the latent domain, perform all subsequent operations in this domain, and decode the result with the decoder.
The bit rate of this approach is determined by the size of the codebooks $K$, and the number of sampling timesteps $T$. Specifically, the bit-stream length is given by $(T-1)\log_{2}(K)$. Therefore, the bit rate can be reduced by simply decreasing the number of codebooks, or by using a smaller number of timesteps at generation, e.g., by skipping every other step, or by using specific timestep intervals (see LABEL:app:range_t). In the approach described so far, the length of the bitstream increases logarithmically with $K$, making it computationally demanding to increase the bit rate. For instance, even for $K=8192$, $T=1000$ and $768\times 768$ images our BPP is approximately 0.022. Thus, to produce higher bit rates, we propose to refine the noise selected at timestep $i$ by employing matching pursuit (MP) (Mallat & Zhang, 1993). Specifically, at each step $i$, we construct the chosen noise as a convex combination of $M$ elements from ${\mathcal{C}}_{i}$, gathered in a greedy fashion to best correlate with the guiding residual ${\mathbf{x}}_{0}-\hat{{\mathbf{x}}}_{0|i}$ (as in Eq. 7). The resulting convex combination involves $M-1$ quantized scalar coefficients, chosen from a finite set of $C$ values taken from $[0,1]$. Therefore, the resulting length of the bit-stream is given by $\smash{(T-1)(\log_{2}(K)M+C(M-1))}$, such that $M=1$ is similar to our standard compression scheme, and the length of the bit-stream increases linearly with $M$ and $C$. We apply this algorithm when the absolute bits number crosses $(T-1)\cdot\log_{2}(2^{13})$. Further details are available in LABEL:app:matching_pursuit.
Experiments.
We evaluate our compression method on Kodak24 (Franzen, 1999), DIV2K (Agustsson & Timofte, 2017), ImageNet 1K $256\times 256$ (Deng et al., 2009; Pan et al., 2020), and CLIC2020 (Toderici et al., 2020). For all datasets but ImageNet, we center crop and resize all images to $512\times 512$. We compare to numerous competing methods, both non-neural and neural, and both zero-shot, fine-tuning based, and training based approaches. For the ImageNet dataset, we use the unconditional pixel space ImageNet $256\times 256$ model of Dhariwal & Nichol (2021), and compare our results to BPG (Bellard, 2018), HiFiC (Mentzer et al., 2020), IPIC (Xu et al., 2024), and two PSC (Elata et al., 2024) configurations, distortion-oriented (PSC-D) and perception-oriented (PSC-P). For all other datasets, we use SD 2.1 $512\times 512$ (Rombach et al., 2022) and compare to BPG, PSC-D, PSC-P, ILLM (Muckley et al., 2023), PerCo (SD) (Körber et al., 2024; Careil et al., 2023), and twoCRDR (Iwai et al., 2024) configurations, distortion-oriented (CRDR-R) and perception-oriented (CRDR-R). PSC shares the same pre-trained model as ours, while PerCo (SD) requires additional fine-tuning. For our method, we apply SD 2.1 unconditionally, as we saw no improvement by adding prompts (see further details in LABEL:app:text_effect). We assess our method for several options of $T$, $K$, $M$, and $C$ to control the bit rate. See further details in LABEL:app:compression_details. We evaluate distortion with PSNR and LPIPS (Zhang et al., 2018) and perceptual quality with FID (Bińkowski et al., 2018). For ImageNet, FID is computed against the entire 50k $256\times 256$ validation set. For the smaller datasets we follow Mentzer et al. (2020) and compute the FID over extracted image patches. Specifically, for DIV2K and CLIC2020 we extract $128\times 128$ sized patches, and for Kodak we use $64\times 64$.
As shown in Fig. 5, our compression scheme dominates previous methods on the rate-perception-distortion plane (Blau & Michaeli, 2019) for lower bit rates, surpassing both the perceptual quality (FID) and distortion (PSNR and LPIPS) of previous methods. For instance, our FID scores are lower than those of all other methods at around 0.1 BPP, while, for the same BPP, our distortion performance is better than the perceptually-oriented methods (e.g., PerCo, PSC-P, and IPIC). However, our method under-performs at the highest bit rates, especially when using SD. This is expected due to the performance ceiling entailed by the pre-trained encoder-decoder of SD (Körber et al., 2024; Elata et al., 2024). The qualitative comparisons in Fig. 4 further demonstrate our superior perceptual quality, where, even for extreme bit rates, our method maintains similarity to the original images in fine details. See LABEL:app:compression_details for more details and results.
Figure 5: Compression quantitative evaluation. We compare the perceptual quality (FID) and distortion (PSNR, LPIPS) achieved at different BPPs. The image sizes of each dataset is denoted next to its name. Our method produces the best perceptual quality at most BPPs. Importantly, this is while we also attain lower distortion compared to perceptually-oriented methods (e.g., PSC-P and PerCo (SD)). For the three rightmost datasets, note that our approach, PSC-P, and PerCo (SD) use the latent space Stable Diffusion 2.1 model, so its VAE imposes a distortion bound. Thus, we report the distortion attained by simply passing the images through this VAE (dashed line).
6 Compressed Conditional Generation
We showed that DDCM can be used as an image codec by following a simple index selection rule, guiding the generated image towards a target one. Here, we generalize this scheme to any conditional generation task, considering the more broad framework of compressing conditionally generated samples. This is a particularly valuable framework in scenarios where the input condition ${\mathbf{y}}$ is bit rate intensive, e.g., where ${\mathbf{y}}$ is a degraded image and the goal is to produce a compressed high-quality reconstruction of it. To the best of our knowledge, this task, which we name compressed conditional generation, has only been thoroughly explored for text input conditions (Rajesh et al., 2023).
A naive solution to this task could be to simply compress the outputs of any existing conditional generation scheme. Here we propose a novel end-to-end solution that generates the outputs directly in their compressed form. Importantly, note that our approach in Sec. 4 requires the condition ${\mathbf{y}}$ for decompressing the bit-stream. While this is not a stringent requirement when the condition is lightweight (e.g., a text prompt), this approach is less suitable when storage of the condition signal itself requires a long bit-stream. The solutions we propose in this section enable decoding the bit-stream without access to ${\mathbf{y}}$.
Compressed Conditional Generation with DDCM.
We propose generating a conditional sample by choosing the indices $k_{i}$ in Eq. 6 via
$k_{i}=\operatorname*{arg\,min}_{k\in\{1,\ldots,K\}}{\mathcal{L}}({\mathbf{y}},{\mathbf{x}}_{i},{\mathcal{C}}_{i},k),$ (8)
instead of picking them randomly. Here, ${\mathcal{L}}({\mathbf{y}},{\mathbf{x}}_{i},{\mathcal{C}}_{i},k)$ can be any loss function that attains a lower value when ${\mathcal{C}}_{i}(k)$ directs the generative process towards an image that matches ${\mathbf{y}}$. For example, for the loss
${\mathcal{L}}_{\text{P}}({\mathbf{y}},{\mathbf{x}}_{i},{\mathcal{C}}_{i},k)=\lVert{\mathcal{C}}_{i}(k)-\sigma_{i}\nabla_{{\mathbf{x}}_{i}}\log{p_{i}({\mathbf{y}}|{\mathbf{x}}_{i})}\rVert^{2}$ (9)
we obtain the following result (see proof in LABEL:appendix:cond_compression):
Proposition 6.1.
Suppose that image samples are generated via process 6, and the indices $k_{i}$ are chosen according to Eq. 8 with ${\mathcal{L}}={\mathcal{L}}_{\textnormal{P}}$. Then, when $K\rightarrow\infty$, such a generative process becomes a discretization of a probability flow ODE over the posterior distribution $p_{0}({\mathbf{x}}_{0}|{\mathbf{y}})$.
In other words, Proposition 6.1 implies that for the loss ${\mathcal{L}}_{\text{P}}$, increasing $K$ leads to more accurate sampling from the posterior $p_{0}({\mathbf{x}}_{0}|{\mathbf{y}})$, though this results in longer bit-streams. Thus, as long as we have access to $\nabla_{{\mathbf{x}}_{i}}\log{p_{i}({\mathbf{y}}|{\mathbf{x}}_{i})}$ (or an approximation of it) ${\mathcal{L}}_{\text{P}}$ may serve as a sensible option for solving a compressed conditional generation task with DDCM. Interestingly, we show in LABEL:appendix:compression_private_case that our compression scheme from Sec. 5 is a special case of the proposed compressed conditional generation method, with ${\mathbf{y}}={\mathbf{x}}_{0}$ and ${\mathcal{L}}={\mathcal{L}}_{\text{P}}$.
6.1 Compressed Posterior Sampling for Image Restoration
Figure 6: Comparison of zero-shot posterior sampling image restoration methods. Our approach achieves better perceptual quality compared to previous methods, while maintaining competitive PSNR and automatically producing compressed bit-stream representations for each restored image.
Our compressed conditional sampling approach can be utilized for solving inverse problems via posterior sampling. Specifically, we consider inverse problems of the form ${\mathbf{y}}={\bm{A}}{\mathbf{x}}_{0}$, where ${\bm{A}}$ is some linear operator. We restrict our attention to unconditional diffusion models and solve the problems in a “zero-shot” manner (similarly to Kawar et al. (2022); Chung et al. (2023); Wang et al. (2023b)). To generate conditional samples, we propose using the loss
${\mathcal{L}}({\mathbf{y}},{\mathbf{x}}_{i},{\mathcal{C}}_{i},k)=\lVert{\mathbf{y}}-{\bm{A}}({\bm{\mu}}_{i}({\mathbf{x}}_{i})+\sigma_{i}{\mathcal{C}}_{i}(k))\rVert^{2}.$ (10)
Note that Eq. 10 attains a lower value when $\sigma_{i}{\bm{A}}{\mathcal{C}}_{i}(k)$ points in the direction that perturbs ${\bm{A}}{\bm{\mu}}_{i}({\mathbf{x}}_{i})$ towards ${\mathbf{y}}$. Thus, our conditional generative process aims to produce a reconstruction $\hat{{\mathbf{x}}}$ that satisfies ${\bm{A}}\hat{{\mathbf{x}}}\approx{\mathbf{y}}$, implying that we approximate posterior sampling (Ohayon et al., 2023). Notably, when assuming that $p_{i}({\mathbf{y}}|{\mathbf{x}}_{i})$ is a multivariate normal distribution centered around ${\bm{A}}{\mathbf{x}}_{i}$ (as in (Jalal et al., 2021)), the chosen codebook noise ${\mathcal{C}}_{i}(k_{i})$ approximates the gradient $\nabla_{{\mathbf{x}}_{i}}\log{p_{i}({\mathbf{y}}|{\mathbf{x}}_{i})}$ and Eq. 10 becomes a proxy of Eq. 9.
Following (Chung et al., 2023; Wang et al., 2023b), we implement our method using the unconditional ImageNet $256\times 256$ DDM trained by Dhariwal & Nichol (2021). We fix $K=4096$ for all codebooks, resulting in a compressed bit-stream of approximately $0.183$ BPP for each generated image. We compare our method with DPS (Chung et al., 2023) and DDNM (Wang et al., 2023b) on two noiseless tasks: colorization and 4×4$\times$4 × super-resolution (using the bicubic kernel). We evaluate these methods using their official implementations and the same DDM. We additionally compress the outputs of DPS and DDNM to assess whether such a naive approach would yield better results. To do so, we adopt our proposed compression scheme (from Sec. 5), employing the same unconditional ImageNet DDM and using $K=4096$ noises per codebook.
Qualitative and quantitative results are reported in Fig. 6. As expected, due to the rate-perception-distortion tradeoff (Blau & Michaeli, 2019), we observe that compressing the outputs of DPS and DDNM harms either their perceptual quality (FID), or their distortion (PSNR), or both. This is while our method achieves superior perceptual quality compared to both DPS and DDNM, including their compressed versions. While our method achieves slightly worse PSNR, this is expected due to the perception-distortion trade-off (Blau & Michaeli, 2018). See LABEL:appendix:zero-shot for more details.
6.2 Compressed Real-World Face Image Restoration
Figure 7: Comparing real-world face image restoration methods on the WIDER-Test dataset. We successfully optimize the NR-IQA measures and produce appealing output perceptual quality with less artifacts compared to previous methods.
Real-world face image restoration is the practical task of restoring any degraded face image, without any knowledge of the corruption process it has gone through (Wang et al., 2021; Gu et al., 2022; Wang et al., 2022; Zhou et al., 2022; Wang et al., 2023c; Lin et al., 2024; Yue & Loy, 2024; Chen et al., 2024b; Ohayon et al., 2024). We propose a novel method capable of optimizing any no-reference image quality assessment (NR-IQA) measure at test time (e.g., NIQE (Mittal et al., 2013)), without relying on gradients.
Specifically, at each timestep $i$, we start by picking two indices – one that promotes high perceptual quality, $k_{i,P}$, and another that promotes low distortion, $k_{i,D}$. Then, we choose between $k_{i,P}$ and $k_{i,D}$ the index that better optimizes a desired balance of the perception-distortion tradeoff (Blau & Michaeli, 2018). Formally, letting ${\bm{r}}({\mathbf{y}})\approx\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{y}}]$ denote the approximate Minimum Mean-Squared-Error (MMSE) estimator of this task, we pick $k_{i,D}$ via
$k_{i,D}=\operatorname*{arg\,max}_{k\in\{1,\ldots,K\}}\langle{\mathcal{C}}_{i}(k),{\bm{r}}({\mathbf{y}})-\hat{{\mathbf{x}}}_{0|i}\rangle.$ (11)
[…]
$k_{i}\!=\!\operatorname*{arg\,min}_{k\in\{k_{i,D},k_{i,P}\}}\!\text{MSE}\!\left(\!{\bm{r}}({\mathbf{y}}),\hat{{\mathbf{x}}}_{0|i-1}^{(k)}\!\right)\!+\!\lambda Q\!\left(\!\hat{{\mathbf{x}}}_{0|i-1}^{(k)}\!\right)\!,$ (12)
where $Q(\cdot)$ can be any NR-IQA measure, even a non-differentiable one. In LABEL:app:dmax-ot we explain our choice to set ${\bm{r}}({\mathbf{y}})$ as an MMSE estimator.
[…]
The results for the WIDER-Test dataset are reported in Fig. 7 (see LABEL:app:dmax-ot for the other datasets). Our approach clearly optimizes each quality measure effectively and generalizes well according to the $\text{FD}_{\text{DINOv2}}$ scores. This is also confirmed visually, where all of our solutions produce high-quality images with less artifacts compared to previous methods. While our approach shows slightly worse distortion, this is once again expected due to the perception-distortion tradeoff (Blau & Michaeli, 2018).
7 Discussion
We introduced DDCM, a novel generative approach for DDMs that produces high-quality image samples jointly with their lossless compressed bit-stream representations. We found that DDCM achieves comparable generative performance to DDPM, even when the codebooks are extremely small. We leveraged DDCM to solve several compressed image generation tasks, including image compression and compressed restoration, where we achieved state-of-the-art results. Besides image restoration, our compressed conditional generation framework can be used for any type of diffusion guidance, e.g., for text-conditional generation. We demonstrate this option in LABEL:appendix:classifier-guidance and LABEL:appendix:classifier-free-guidance, introducing new classifier-based and classier-free guidance methods that do not use ${\mathbf{y}}$ for decompression. Moreover, we present in LABEL:app:editing preliminary results for compressed image editing using DDCM, by decompressing an image using a desired edit text prompt.
While our empirical results are encouraging, our work does not explain theoretically why DDCM sampling and our simple index selection strategies work so effectively. We encourage future works to investigate the principles behind the success of our methods. Moreover, when operating in latent space, our codec’s performance is bounded by the underlying VAE, particularly at higher bit rates. Our results at higher bit rates could also be improved through better approaches than our current matching pursuit inspired solution. Additionally, our compression efficiency could be improved through entropy coding of the selected indices, potentially reducing bit rates without sacrificing quality. Lastly, all DDCM-based solutions can be improved further by optimizing the codebooks, e.g., through dictionary learning. To conclude, our work demonstrates promising results across numerous tasks while providing opportunities for both theoretical analysis and practical improvements.
Acknowledgements
This research was partially supported by the Israel Science Foundation (ISF) under Grants 2318/22, 951/24 and 409/24, and by the Council for Higher Education – Planning and Budgeting Committee. We thank Noam Elata and Matan Kleiner for assisting with our figures and experiments.